OpenStudy (anonymous):

Li ran 5/6 of a mile from her house, turned around, and ran 2/3 of a mile back. At that point, how far was she from her house? 1/3 mi 1 2/3 mi 1/6 mi 1/2 mi

OpenStudy (anonymous):

5/6 - 2/3 = 5/6 - 4/6 = 1/6 mi

OpenStudy (anonymous):

5/6 - 2/3 = 1/6. So, 1/6 is the answer